OpenStudy (anonymous):

a trail mix company has 200 lb of peanuts and 160 lb of raisins. a batch oh house mix takes 20 lb of peanuts and 10 lb of raisins. a batch of fancy mix takes 10 lb of peanuts and 20 lb of raisins. profit from a batch of house mix is $50 and from a batch of fancy mix is $60. how much of each type should be made to maximize profits? what is the maximum profit?

OpenStudy (anonymous):

lbspeanuts=20h+10f lbsraisins=10h+20f h=batches of house, f=batches of fancy so we have the inequalities 20h+10f <=200 10h+20f<=160 the profit (p): p=50h+60f so we graph the two inequalities, and find where the profit is maximized.

OpenStudy (anonymous):

so how much of each type should be made to maximize the profits? and what is the max profit?

OpenStudy (anonymous):

Look at the boundary, and specifically the endpoints in this case, and I think the maximum is where the boundary lines intersect at h=8, f=4 which gives a profit of: 50*8+60*4=400+240=$640

OpenStudy (anonymous):

THANK YOU!<3

OpenStudy (anonymous):

thank youuuuuuu!